OpenStudy (anonymous):

Katie wants to hang a painting in a gallery but the painting and frame must have an area of 45 square feet. The painting is 6 feet wide by 7 feet long. The frame must have an equal width of x on each side. Which quadratic equation represents the area of the painting and frame combined?

OpenStudy (anonymous):

Do u have options?

OpenStudy (anonymous):

yeah hold on

OpenStudy (anonymous):

4x2 + 26x + 45 = 0 4x2 + 26x − 3 = 0 x2 + 13x − 3 = 0 x2 + 13x + 45 = 0

OpenStudy (anonymous):

Ok...nevermind! i got it! the equation without factoring is (2x + 7)(2x +6) based on the information given! now just foil it out! (2x + 7)(2x +6) = 0 4x^2 + 12x + 14 X + 42 4x^2 + 26X + 42 = 0! Enjoy!

OpenStudy (anonymous):

oh wait a second! small error!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

(2x + 7)(2x + 6) = 45 Not 0! then, it'll be 4x^2 + 26x + 42 - 45 = 0 then, the final answer is 4x^2 + 26x - 3 = 0! So, the answer is B!

OpenStudy (anonymous):

Can u understand?

OpenStudy (anonymous):

yeah i get it now! thank you! (: